# Deep Positive-Unlabeled Anomaly Detection for Contaminated Unlabeled Data

### v1

@@ 3.2 Application to Other Anomaly Detectors
The PUAE is based on the loss function of the AE defined in Eq. (1), and can be applied to other anomaly detectors by substituting their loss function into Eq. (1). This application requires that their loss function be non-negative and differentiable. A lot of approaches satisfy this requirement, such as the denoising AE (DAE) [55] and deep support vector data description (DeepSVDD) [46].
One example is the DAE, which is a variant of the AE, and has also been successfully applied to anomaly detection [50]. The DAE aims to reconstruct original data points from noisy input data points. Its loss function for each data point is defined as follows:
@@
where $\epsilon$ is a noise from an isotropic Gaussian distribution. Our approach can be applied to the DAE by substituting this into Eq. (1).
The other example is the DeepSVDD, which aims to pull the representation of the normal data towards the predefined center, and push those of the anomaly data away from it. Let $f_{\theta}(\mathbf{x})$ be the feature extractor, like the encoder in the AE. The loss function for each data point of the DeepSVDD is defined as follows:
$\ell(\mathbf{x};\theta)=\left\|f_{\theta}(\mathbf{x})-\mathbf{c}\right\|^{2},$ (15)
where $\mathbf{c}\neq\mathbf{0}$ is the predefined center vector. Our approach can be applied to the DeepSVDD by substituting this into Eq. (1).
@@ 4.1 Unsupervised Anomaly Detection
Numerous unsupervised anomaly detectors have been presented, ranging from shallow approaches such as the one-class support vector machine (OCSVM) [54] and the isolation forest (IF) [27] to deep approaches such as the AE [16] and the DeepSVDD [46]. In addition, generative models such as the variational autoencoder [20, 21] and the generative adversarial nets [12] are also used for anomaly detection [5, 52, 45, 40, 58, 14, 60]. Although they are often used in anomaly detection, they have limitations in detection performance because they cannot use information about anomalies. For example, generative models may fail to detect anomalies that are obvious to the human eye [32].
Furthermore, these approaches assume that unlabeled data are mostly normal. However, they are contaminated with anomalies in practice, degrading the detection performance. Several unsupervised approaches have been presented to handle such contaminated unlabeled data [63, 43, 53]. Among them, the latent outlier exposure (LOE) [43] is one of the most representative methods. The LOE introduces the label for each data point as the latent variable, and alternates between inferring the latent label and optimizing the parameter of the base anomaly detector. Compared to these unsupervised approaches, our approach can achieve high anomaly detection performance by using the unlabeled and anomaly data, even if the unlabeled data are contaminated with anomalies. As the base anomaly detector for our approach, the AE and the DeepSVDD can be used as described in Section 3. In addition, our approach can also be applied to the LOE by substituting its objective function into $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$ in Eq. (11).
@@ 4.2 Semi-supervised Anomaly Detection
Several semi-supervised anomaly detectors, aiming to improve the detection performance using anomaly data in addition to unlabeled data, have been presented [15, 59, 47]. Among them, the DeepSAD [47] is one of the most representative methods. The DeepSAD trains the DeepSVDD model to minimize Eq. (15) for the unlabeled data, and to maximize it for the anomaly data. Compared to these semi-supervised approaches, our approach can handle the unlabeled data that are contaminated with anomalies, as described in Section 3.1.
To handle contaminated unlabeled data, several semi-supervised approaches have been presented, including PU learning approaches [62, 18, 61, 37, 25, 41]. Among them, the semi-supervised outlier exposure with a limited labeling budget (SOEL) [25] is the current state-of-the-art approach. The SOEL is a semi-supervised extension of the LOE [43], and presents the query strategy for the LOE, deciding which data should be labeled. We compare our approach with the SOEL in Section 5, and confirm the effectiveness of our approach.
@@ 4.3 Positive-Unlabeled Learning
A lot of PU learning approaches have been presented in the context of binary classification [10, 9, 8, 22, 2, 31]. Among them, our approach is based on the unbiased PU learning [9, 8, 22]. The empirical risk estimators in Eq. (11) is unbiased and consistent with respect to all popular loss function. This means for fixed $\theta$, Eq. (11), which is the approximation of Eq. (8), converges to Eq. (8) as the dataset sizes $N,M\rightarrow\infty$. It is known that if the model is linear with respect to $\theta$, a particular loss function will result in convex optimization, and the globally optimal solution can be obtained [33, 38, 35]. Despite this ideal property, when the model is complex such as neural networks, Eq. (11) can be negative, potentially leading to the meaningless solution. To avoid this, according to [22], we use the max operator to prevent it from being negative.
Compared to conventional PU learning that is based on the binary classifier, our approach is based on the anomaly detector such as the AE. Although conventional PU learning cannot detect unseen anomalies since its decision boundary is between normal data points and seen anomalies, our approach can detect both seen and unseen anomalies.
@@ 5.1 Data
We used following image datasets: MNIST [51], FashionMNIST [57], SVHN [34], and CIFAR10 [23]. MNIST is the handwritten digits, FashionMNIST is the fashion product images, SVHN is the house number digits, and CIFAR10 is the animal and vehicle images. We converted SVHN and CIFAR10 into gray-scale, and resized all datasets to $32\times 32$ resolution.
These datasets consist of 10 class images. For MNIST and SVHN, we used the digits as class indices. For FashionMNIST, we indexed labels as: {T-shirt/top: 0, Trouser: 1, Pullover: 2, Dress: 3, Coat: 4, Sandal: 5, Shirt: 6, Sneaker: 7, Bag: 8, and Ankle boot: 9}. For CIFAR10, we indexed labels as: {airplane: 0, automobile: 1, bird: 2, cat: 3, deer: 4, dog: 5, frog: 6, horse: 7, ship: 8, and truck: 9}.
We extend the experiments in [46] to semi-supervised anomaly detection with contaminated unlabeled data. Of the 10 classes, we used one class as normal, another class as unseen anomaly, and the remaining classes as seen anomaly. For example with MNIST, if we use the digit 1 as normal and the digit 0 as unseen anomaly, seen anomaly corresponds to the digits 2, 3, 4, 5, 6, 7, 8, and 9. For all datasets, we used class 0 as unseen anomaly, and select one normal class from the remaining 9 classes. The training dataset consists of 5,000 samples, of which 4,500 samples are unlabeled normal data points, 250 samples are labeled seen anomalies, and 250 samples are unlabeled seen anomalies. That is, the unlabeled data points in this dataset are contaminated with seen anomalies. We used 10% of the training dataset as the validation dataset. The test dataset consists of 2,000 samples, about half of which are normal and the rest are anomalies, including both seen and unseen anomalies. More precisely, the normal data points are sampled from the test dataset of the normal class up to 1,000 samples, and the rest are sampled from the other classes. We provide the example of the dataset in the case of MNIST in Appendix A.
@@ 5.2 Methods
We applied our approach to the AE [16] and the DeepSVDD [46]. We refer to the former as PUAE, and the latter as PUSVDD. We compared our PUAE and PUSVDD with the following unsupervised and semi-supervised approaches.
Unsupervised approaches: We used the IF [27] as the shallow approach, and used the AE [16] and the DeepSVDD [46] as the deep approaches. We also used the LOE [43], which is robust to the anomalies in the unlabeled data. We chose the AE as the base detector for the LOE.
Semi-supervised approaches: We used the ABC [59], the DeepSAD [47], and the SOEL [25] that is a semi-supervised extension of the LOE. We chose the AE as the base detector for the SOEL. We also used the PU learning binary classifier (PU) [22] for reference.
@@ 5.3 Setup
First, we outline the setups for deep approaches, which include all approaches except the IF. We used two-hidden-layer neural networks (1,000 units per layer) for the encoder and the decoder in the AE, the feature extractor in the DeepSVDD, and the binary classifier in the PU. We used the ReLU as the activations for the above neural networks. For the AE-based and DeepSVDD-based approaches, we set the dimension of the latent variable to 40. For the DeepSVDD-based approaches, we used no bias terms in each layer, pretrained these feature extractors as the AE, and set the center $\mathbf{c}$ in Eq. (15) to the mean of the outputs of the encoder. We trained all methods by using Adam [19] with a mini-batch size of 128. We set the learning rate to $10^{-4}$, set the maximum number of epochs to 200, used the weight decay [13] with $10^{-3}$, and used early-stopping [13] based on the validation dataset. We set the hyperparameter $\alpha=0.1$ for the PUAE, the PUSVDD, the PU, the LOE, and the SOEL, which represents the probability of anomaly occurrence.
We next outline the setup for the IF. We used the implementation provided by scikit-learn [39]. The hyperparameters were left as default in our experiments.
For unsupervised approaches, we trained them only using the unlabeled data in the training dataset. Note that we did NOT use the anomaly data since unsupervised approaches cannot effectively use them. For semi-supervised approaches, we trained them using the unlabeled and anomaly data.
To measure the anomaly detection performance, we calculated the AUROC scores for all datasets. We ran all experiments eight times while changing the random seeds.
The machine specifications used in the experiments are as follows: the CPU is Intel Xeon Platinum 8360Y, the memory size is 512GB, and the GPU is NVIDIA A100 SXM4.
@@ 5.4 Results
Table 1: Comparison of anomaly detection performance on image datasets. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value.
@@
IF [27] 0.829 $\pm$ 0.089 0.911 $\pm$ 0.060 0.514 $\pm$ 0.013 0.561 $\pm$ 0.097
AE [16] 0.894 $\pm$ 0.056 0.823 $\pm$ 0.094 0.598 $\pm$ 0.027 0.532 $\pm$ 0.127
DeepSVDD [46] 0.763 $\pm$ 0.060 0.667 $\pm$ 0.070 0.513 $\pm$ 0.014 0.556 $\pm$ 0.032
LOE [43] 0.905 $\pm$ 0.050 0.860 $\pm$ 0.101 0.594 $\pm$ 0.028 0.533 $\pm$ 0.125
ABC [59] 0.897 $\pm$ 0.055 0.826 $\pm$ 0.093 0.599 $\pm$ 0.027 0.533 $\pm$ 0.127
DeepSAD [47] 0.795 $\pm$ 0.054 0.693 $\pm$ 0.072 0.514 $\pm$ 0.016 0.560 $\pm$ 0.033
SOEL [25] 0.944 $\pm$ 0.034 0.896 $\pm$ 0.075 0.601 $\pm$ 0.032 0.535 $\pm$ 0.123
PU [22] 0.943 $\pm$ 0.071 0.811 $\pm$ 0.180 0.500 $\pm$ 0.014 0.488 $\pm$ 0.066
PUAE 0.971 $\pm$ 0.019 0.940 $\pm$ 0.047 0.608 $\pm$ 0.030 0.554 $\pm$ 0.115
PUSVDD 0.922 $\pm$ 0.029 0.884 $\pm$ 0.060 0.538 $\pm$ 0.030 0.575 $\pm$ 0.037
Table 1 compares the anomaly detection performance on each image dataset. First, we focus on the unsupervised approaches. The IF showed the high detection performance on FashionMNIST and CIFAR10. This indicates that the tree structure is somewhat effective on the contaminated datasets. The deep approaches such as the AE and the DeepSVDD are more easily influenced by the contaminated datasets due to their flexibility. As a result, they performed worse than our PUAE and PUSVDD in all cases. The DeepSVDD performed worse than the AE except for CIFAR10. This indicates that the DeepSVDD is more easily influenced by the contaminated datasets than the AE. One possible reason is that the training objective function of the DeepSVDD (Eq. (15)) is much simpler than that of the AE (Eq. (1)). Although the LOE does not use the anomaly data, it can mitigate the effect of the contaminated datasets to some extent. In particular, it significantly improved the detection performance in FashionMNIST compared to the AE, which is the base detector of the LOE.
Next, we focus on semi-supervised approaches. The ABC and the DeepSAD improved the detection performance compared to the AE and the DeepSVDD, respectively. This indicates that we can improve the detection performance by using the anomaly data. However, their improvements were marginal because the unlabeled data are contaminated with anomalies. The SOEL, which can handle such unlabeled data, achieved the second best performance compared to our approaches. These results indicate the effectiveness of our approaches in handling the contaminated unlabeled data.
Finally, we focus on our approaches. In almost all cases, our PUAE and PUSVDD performed better than the AE and the ABC, and the DeepSVDD and the DeepSAD, respectively. Especially, in MNIST, FashionMNIST, and SVHN, the PUAE achieved the best performance among all detectors. This shows the effectiveness of our approach, which is based on PU learning and the anomaly detector. Although the PUAE performed poorly in CIFAR10, the PUSVDD achieved the best performance in CIFAR10. This indicates that we can achieve higher detection performance by selecting the base anomaly detector that is better suited to the target dataset.
@@
In addition, we focus on the difference between our approaches and the PU. Although the PU achieved high performance on MNIST and FashionMNIST, its performance was worse than the PUAE in all datasets. The reason is as follows. Since the PU is for binary classification, it sets the decision boundary between normal data points and seen anomalies. This prevents us from detecting unseen anomalies. On the other hand, our approach can detect unseen anomalies since it can model normal data points by the anomaly detector. To support this claim, we provide the anomaly detection performance for seen and unseen anomalies on each dataset in Appendix B.

### v2

@@ 3.3 Application to Other Anomaly Detectors
Our framework is applicable to anomaly detection models whose loss functions are non-negative and differentiable. As described above, the AE and the DeepSVDD satisfy this requirement. In addition, a lot of models satisfy this requirement, such as the DAE (Vincent et al., 2008) and recent self-supervised detectors (Hendrycks et al., 2019; Qiu et al., 2021; Shenkar and Wolf, 2021). For example, we focus on the DAE. The DAE is a variant of the AE, and has also been successfully applied to anomaly detection (Sakurada and Yairi, 2014). The DAE aims to reconstruct original data points from noisy input data points. Its loss function for each data point is defined as follows:
@@
where $\epsilon$ is a noise from an isotropic Gaussian distribution. Our framework can be applied to the DAE by substituting this into Eq. (1). In this way, our framework can be applied by substituting the loss function of the desired model into Eq. (1) in the PUAE or Eq. (14) in the PUSVDD.
@@ 4.1 Unsupervised Anomaly Detection
Numerous unsupervised approaches have been presented, ranging from shallow approaches such as the one-class support vector machine (OCSVM) (Tax and Duin, 2004) and the isolation forest (IF) (Liu et al., 2008) to deep approaches such as the AE (Hinton and Salakhutdinov, 2006) and the DeepSVDD (Ruff et al., 2018). In addition, generative models such as the variational autoencoder (Kingma and Welling, 2014; Kingma et al., 2015) and the generative adversarial nets (Goodfellow et al., 2014) are also used for anomaly detection (Choi et al., 2018; Serrà et al., 2019; Ren et al., 2019; Perera et al., 2019; Xiao et al., 2020; Havtorn et al., 2021; Yoon et al., 2021). Although they are often used in anomaly detection, their detection performance is limited because they cannot use information about anomalies. For example, generative models may fail to detect anomalies that are obvious to the human eye (Nalisnick et al., 2018).
Furthermore, these approaches assume that unlabeled data are mostly normal. However, they are contaminated with anomalies in practice, degrading the detection performance. Several unsupervised approaches have been presented to handle such contaminated unlabeled data (Zhou and Paffenroth, 2017; Qiu et al., 2022; Shang et al., 2023). Among them, the latent outlier exposure (LOE) (Qiu et al., 2022) is a representative approach. The LOE introduces the label for each data point as the latent variable, and alternates between inferring the latent label and optimizing the parameter of the base anomaly detector. Compared to these approaches, our approach can achieve better detection performance by using the unlabeled data and the labeled anomaly data, even if the unlabeled data are contaminated with anomalies. As the base detector for our approach, the AE, the DeepSVDD, and recent self-supervised detectors (Hendrycks et al., 2019; Qiu et al., 2021; Shenkar and Wolf, 2021) can be used as described in Section 3. In addition, our approach can also be applied to the LOE by substituting its objective function into $\mathcal{L}_{\mathcal{U}}^{-}(\theta)$ in Eq. (11).
@@ 4.2 Semi-supervised Anomaly Detection
Several semi-supervised approaches have been presented, aiming to improve the detection performance using labeled anomaly data in addition to unlabeled data (Hendrycks et al., 2018; Yamanaka et al., 2019; Ruff et al., 2019). Compared to these approaches, our approach can effectively handle the unlabeled data that are contaminated with anomalies, as described in Section 3.1.
To handle contaminated unlabeled data, a number of semi-supervised approaches have been presented, including PU learning approaches (Zhang et al., 2018; Ju et al., 2020; Zhang et al., 2021; Pang et al., 2023; Li et al., 2023; Perini et al., 2023). Among them, the semi-supervised outlier exposure with a limited labeling budget (SOEL) (Li et al., 2023) is the current state-of-the-art approach. The SOEL is a semi-supervised extension of the LOE (Qiu et al., 2022), and presents the query strategy for the LOE, deciding which data should be labeled. Compared to these approaches, our approach is theoretically justified from the perspective of unbiased PU learning (Du Plessis et al., 2014, 2015; Kiryo et al., 2017). In addition, our approach achieved equal to or better performance than the SOEL across eight datasets, as described in Section 5.
@@ 4.3 Positive-Unlabeled Learning
A lot of PU learning approaches have been presented for binary classification (Elkan and Noto, 2008; Du Plessis et al., 2014, 2015; Kiryo et al., 2017; Bekker and Davis, 2020; Nakajima and Sugiyama, 2023). Among them, our approach is based on the unbiased PU learning (Du Plessis et al., 2014, 2015; Kiryo et al., 2017). The empirical risk estimators in Eq. (11) is unbiased and consistent with respect to all popular loss function. This means for fixed $\theta$, Eq. (11), which is the approximation of Eq. (8), converges to Eq. (8) as the dataset sizes $N,M\rightarrow\infty$. It is known that if the model is linear with respect to $\theta$, a particular loss function will result in convex optimization, and the globally optimal solution can be obtained (Natarajan et al., 2013; Patrini et al., 2016; Niu et al., 2016). Despite this ideal property, when the model is complex such as neural networks, Eq. (11) can become negative, potentially leading to the meaningless solution. To address this issue, following (Hammoudeh and Lowd, 2020), we take its absolute value, as described in Section 3.1.
Compared to conventional PU learning that is based on the binary classifier, our approach is based on deep anomaly detection models such as the AE and the DeepSVDD. Although conventional PU learning cannot detect unseen anomalies since its decision boundary is between normal data points and seen anomalies, our approach can detect both seen and unseen anomalies.
@@ 5.1 Data
We used following eight image datasets: MNIST (Salakhutdinov and Murray, 2008), FashionMNIST (Xiao et al., 2017), SVHN (Netzer et al., 2011), CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky et al., 2009), PathMNIST, OCTMNIST, and TissueMNIST (Yang et al., 2021, 2023).
First, we explain the first four datasets. MNIST is the handwritten digits, FashionMNIST is the fashion product images, SVHN is the house number digits, and CIFAR10 is the animal and vehicle images. We resized all datasets to $32\times 32$ resolution. These datasets consist of 10 class images. For MNIST and SVHN, we used the digits as class indices. For FashionMNIST, we indexed labels as: {T-shirt/top: 0, Trouser: 1, Pullover: 2, Dress: 3, Coat: 4, Sandal: 5, Shirt: 6, Sneaker: 7, Bag: 8, and Ankle boot: 9}. For CIFAR10, we indexed labels as: {airplane: 0, automobile: 1, bird: 2, cat: 3, deer: 4, dog: 5, frog: 6, horse: 7, ship: 8, and truck: 9}. We extend the experiments in (Ruff et al., 2018) to semi-supervised anomaly detection with contaminated unlabeled data. Of the 10 classes, we used one class as normal, another class as unseen anomaly, and the remaining classes as seen anomaly. For example with MNIST, if we use the digit 1 as normal and the digit 0 as unseen anomaly, seen anomaly corresponds to the digits 2, 3, 4, 5, 6, 7, 8, and 9. For all datasets, we used class 0 as unseen anomaly, and select one normal class from the remaining 9 classes. The training dataset consists of 5,000 samples, of which 4,500 samples are unlabeled normal data points, 250 samples are labeled seen anomalies, and 250 samples are unlabeled seen anomalies. That is, the unlabeled data points in this dataset are contaminated with seen anomalies. We used 10% of the training dataset as the validation dataset. The test dataset consists of 2,000 samples, about half of which are normal and the rest are anomalies, including both seen and unseen anomalies. More specifically, normal data points are sampled from the normal class in the test dataset, with a maximum of 1,000 samples, while both seen and unseen anomalies are sampled from their respective classes, with a maximum of 500 samples each. The example of the MNIST dataset is provided in Appendix A.
Next, we explain the last four datasets. CIFAR100 is just like CIFAR10 but consists of 100 classes. These classes are grouped into 20 superclasses, from which we used nature-related classes as normal and human-related classes as anomalies. We used the people class as unseen anomaly. PathMNIST, OCTMNIST, and TissueMNIST are medical image datasets. PathMNIST is colorectal cancer histology dataset with 9 tissue types. OCTMNIST is retinal optical coherence tomography dataset with 4 diagnostic categories. TissueMNIST kidney cortex cell dataset with 8 categories. For PathMNIST and TissueMNIST, we used the first class as unseen anomaly, selected one class from the remaining ones as normal, and treated the rest as seen anomaly. For OCTMNIST, since a pre-defined normal class exists, we used it as normal, used the first class as unseen anomaly, and treated the remaining classes as seen anomaly.
@@ 5.2 Methods
We compared our PUAE and PUSVDD with the following unsupervised and semi-supervised approaches.
Unsupervised approaches: We used the IF (Liu et al., 2008) as the shallow approach, and used the AE (Hinton and Salakhutdinov, 2006) and the DeepSVDD (Ruff et al., 2018) as the deep approaches. We also used the LOE (Qiu et al., 2022), which is robust to the anomalies in the unlabeled data. We chose the DeepSVDD as the base detector for the LOE.
Semi-supervised approaches: We used the ABC (Yamanaka et al., 2019), the DeepSAD (Ruff et al., 2019), and the SOEL (Li et al., 2023) that is a semi-supervised extension of the LOE. We chose the DeepSVDD as the base detector for the SOEL. We also used the PU learning binary classifier (PU) (Kiryo et al., 2017) for reference.
@@ 5.3 Setup
First, we outline the setups for all approaches except the IF. We used convolutional neural networks for the encoder and the decoder in the AE, the feature extractor in the DeepSVDD, and the binary classifier in the PU. The network architecture follows (Ruff et al., 2018). For the AE-based and DeepSVDD-based approaches, we set the dimension of the latent variable to 128. For the DeepSVDD-based approaches, we used no bias terms in each layer, pre-trained these feature extractors as the AE, and set the center $\mathbf{c}$ in Eq. (14) to the mean of the outputs of the encoder. We trained all methods by using Adam (Kingma and Ba, 2015) with a mini-batch size of 128. We set the learning rate to $10^{-4}$ and the maximum number of epochs to 200. We also used the weight decay (Goodfellow et al., 2016) with $10^{-3}$ and used early-stopping (Goodfellow et al., 2016) based on the validation dataset. We set the hyperparameter $\alpha=0.1$ for the PUAE, the PUSVDD, the PU, the LOE, and the SOEL, which is the probability of anomaly occurrence.
Next, we outline the setup for the IF. We used the scikit-learn implementation (Pedregosa et al., 2011) and kept all hyperparameters at their default values in our experiments.
We trained unsupervised approaches using the unlabeled data222 Note that we did NOT use the labeled anomaly data since unsupervised approaches cannot effectively use them. , while we trained semi-supervised approaches using the unlabeled data and the labeled anomaly data.
To measure the detection performance, we calculated the AUROC scores for all datasets. We ran all experiments five times while changing the random seeds.
The machine specifications used in the experiments are as follows: the CPU is AMD EPYC 9124 16-Core Processor, the memory size is 512GB, and the GPU is NVIDIA RTX 6000 Ada.
@@ 5.4 Results
Table 1: Comparison of anomaly detection performance on the first four datasets.
@@
IF (Liu et al., 2008) 0.885 $\pm$ 0.062 0.916 $\pm$ 0.077 0.501 $\pm$ 0.014 0.610 $\pm$ 0.095
AE (Hinton and Salakhutdinov, 2006) 0.912 $\pm$ 0.042 0.841 $\pm$ 0.102 0.562 $\pm$ 0.040 0.535 $\pm$ 0.120
DeepSVDD (Ruff et al., 2018) 0.937 $\pm$ 0.045 0.921 $\pm$ 0.088 0.582 $\pm$ 0.035 0.709 $\pm$ 0.054
LOE (Qiu et al., 2022) 0.945 $\pm$ 0.033 0.916 $\pm$ 0.094 0.624 $\pm$ 0.054 0.718 $\pm$ 0.062
ABC (Yamanaka et al., 2019) 0.916 $\pm$ 0.042 0.841 $\pm$ 0.104 0.562 $\pm$ 0.041 0.535 $\pm$ 0.119
DeepSAD (Ruff et al., 2019) 0.942 $\pm$ 0.041 0.928 $\pm$ 0.089 0.652 $\pm$ 0.034 0.726 $\pm$ 0.051
SOEL (Li et al., 2023) 0.965 $\pm$ 0.026 0.936 $\pm$ 0.079 0.727 $\pm$ 0.043 0.775 $\pm$ 0.057
PU (Kiryo et al., 2017) 0.962 $\pm$ 0.034 0.922 $\pm$ 0.092 0.681 $\pm$ 0.096 0.693 $\pm$ 0.120
PUAE 0.983 $\pm$ 0.015 0.918 $\pm$ 0.085 0.689 $\pm$ 0.060 0.667 $\pm$ 0.066
PUSVDD 0.989 $\pm$ 0.012 0.948 $\pm$ 0.079 0.747 $\pm$ 0.080 0.803 $\pm$ 0.046
Table 2: Comparison of anomaly detection performance on the last four datasets.
CIFAR100 PathMNIST OCTMNIST TissueMNIST
IF (Liu et al., 2008) 0.604 $\pm$ 0.004 0.809 $\pm$ 0.118 0.714 $\pm$ 0.004 0.472 $\pm$ 0.187
AE (Hinton and Salakhutdinov, 2006) 0.589 $\pm$ 0.010 0.605 $\pm$ 0.240 0.860 $\pm$ 0.005 0.468 $\pm$ 0.178
DeepSVDD (Ruff et al., 2018) 0.587 $\pm$ 0.026 0.759 $\pm$ 0.148 0.726 $\pm$ 0.052 0.661 $\pm$ 0.055
LOE (Qiu et al., 2022) 0.576 $\pm$ 0.035 0.721 $\pm$ 0.160 0.783 $\pm$ 0.030 0.635 $\pm$ 0.086
ABC (Yamanaka et al., 2019) 0.590 $\pm$ 0.010 0.604 $\pm$ 0.241 0.857 $\pm$ 0.001 0.472 $\pm$ 0.177
DeepSAD (Ruff et al., 2019) 0.594 $\pm$ 0.012 0.763 $\pm$ 0.187 0.823 $\pm$ 0.038 0.683 $\pm$ 0.053
SOEL (Li et al., 2023) 0.633 $\pm$ 0.013 0.791 $\pm$ 0.145 0.856 $\pm$ 0.016 0.703 $\pm$ 0.062
PU (Kiryo et al., 2017) 0.541 $\pm$ 0.025 0.807 $\pm$ 0.132 0.614 $\pm$ 0.109 0.633 $\pm$ 0.082
PUAE 0.623 $\pm$ 0.014 0.776 $\pm$ 0.168 0.847 $\pm$ 0.011 0.594 $\pm$ 0.104
PUSVDD 0.637 $\pm$ 0.017 0.831 $\pm$ 0.152 0.857 $\pm$ 0.017 0.731 $\pm$ 0.077
Tables 1 and 2 compare the anomaly detection performance on each dataset. We showed the average of the AUROC scores for all normal classes. We used bold to highlight the best results and statistically non-different results according to a pair-wise $t$-test. We used 5% as the p-value.
First, we focus on unsupervised approaches. The IF and the AE show significant performance variations across different datasets. Although the IF performed well on PathMNIST and the AE performed well on OCTMNIST, their performance became poor on other datasets. The DeepSVDD generally outperformed the IF and the AE, and the LOE, which is based on the DeepSVDD, often performed better than the DeepSVDD. However, since the LOE estimates anomalies within the unlabeled data in an unsupervised manner, incorrect estimations can lead to degraded performance.
Next, we focus on semi-supervised approaches. The ABC and the DeepSAD performed equal to or slightly better than the AE and the DeepSVDD, respectively. The reason for this is that ABC and DeepSAD assume that the unlabeled data are not contaminated with anomalies, which weakens the effect of maximizing anomaly scores for the labeled anomaly data. On the other hand, the SOEL outperformed DeepSAD in all cases. This is because the SOEL is capable of handling anomalies present in the unlabeled data.
Finally, we focus on the proposed methods. In most cases, the PUAE and the PUSVDD performed better than the AE and the ABC, the DeepSVDD and the DeepSAD, respectively. Especially, the PUSVDD achieved the best performance across all datasets. These results strongly indicate the effectiveness of our framework, which integrates PU learning with deep anomaly detection models.
@@
In addition, we focus on the difference between the proposed methods and PU. The proposed methods performed equal to or better than the PU in all datasets. The reason is as follows. Since the PU is for binary classification, it sets the decision boundary between normal data points and seen anomalies. This prevents us from detecting unseen anomalies. On the other hand, our approach can detect unseen anomalies since it can model normal data points by the anomaly detector. The detection performance for seen and unseen anomalies are in Appendix C.